OpenStudy (anonymous):

find all values of x in the interval[0,2pie] that satisfies the equation 2cosx-1=0

OpenStudy (ash2326):

we have \[ 2 \cos x-1=0\] so \[\cos x=\frac{1}{2}\] so \[x= \pi/3, (2\pi-\pi/3)=>x= \pi/3 , 5\pi/3\]

OpenStudy (anonymous):

got it , thank you

OpenStudy (ash2326):

welcome :D